OpenStudy (anonymous):

Need help on differentiating an easy question: y=e^3x+e/x

OpenStudy (anonymous):

I know that when differentiating e^3x, it is 3e^3x But how do you differentiate e/x?

OpenStudy (anonymous):

e/x = ex^(-1)... where e here is considered a constant already... we can use the ff. rule... d [k f(x)] = k d f(x) and for x^(-1), we can use the power formula for differentiation...

OpenStudy (anonymous):

e = 2.718....

OpenStudy (anonymous):

So e/x = e/x^2?

OpenStudy (anonymous):

your exponent is -1...

OpenStudy (anonymous):

ok so final answer is 3e^3x-e/x^2?

OpenStudy (anonymous):

yup...

OpenStudy (anonymous):

One more how do you differentiate 10 sqrt t

OpenStudy (anonymous):



OpenStudy (anonymous):

10t^(1/2)?... we can use power formula also...

OpenStudy (anonymous):

10 as constant...

OpenStudy (anonymous):

So 5t^-1/2

OpenStudy (anonymous):

yup...

OpenStudy (anonymous):

kk thats all thanks

OpenStudy (anonymous):

no problem... :-)